I have a question that states: find the midpoint of segment CD with points (-1,-3) and (3,-1)
Answer by tutorcecilia(2152)

Use the midpoint formula:
Midpoint = (x1+x2)/2 , (y1+y2)/2
Midpoint = x=(-1+3)/2 and y=(-3-1)/2 [plug-in the points (-1,-3) and (3,-1)]
Midpoint = x=2/2 and y= -4/2
Midpoint = (1, -2)